OpenStudy (anonymous):

please show the work onthis problem:30 basketballs (15 Spalding, 10 Wilson, and 5 other brand-name balls) are on a basketball court. Barry Wood closes his eyes and arbitrarily picks up a ball from the court. Determine the probability that the ball selected is a Wilson or a Spalding

OpenStudy (anonymous):

There are 15 chances Barry picks a Spalding, and 10 chances he picks a Wilson, so 25 chances in all out of a total of 30 basketballs. so, probability is 25/30 = 5/6 = 83.33%